OpenStudy (anonymous):

Rewrite the expression without using grouping symbols -6(x + 10) A. (-6) * x + 10 B. (-6) * x + (-6) * 10 C. -6x + 610 D. (-6) + x * (-6) + 10

OpenStudy (anonymous):

does anyone know the answer to this?

OpenStudy (anonymous):

What is the expression?

OpenStudy (anonymous):

the equation above a

OpenStudy (anonymous):

hmm, ok I get it, it needs you to distribute -6 over (x + 10). That means you nee to get rid of the parantheses.

OpenStudy (anonymous):

-6(x=10)

OpenStudy (anonymous):

Let me: \[-6(x + 10) = -6 \times x + (-6) \times 10 = -6x -60\] The answer is D

OpenStudy (anonymous):

thx